OpenStudy (anonymous):

What is the d value in the following arithmetic sequence? -7, -2, 3, 8, .

OpenStudy (texaschic101):

d = 5 -7 + 5 = -2 -2 + 5 = 3 3 + 5 = 8 The difference is 5. So the next number would be : 8 + 5 = 13 then 13 + 5 = 18 then 18 + 5 = 23.....and so on